Circular No. 6095

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

POSSIBLE GRAVITATIONAL MICROLENSING EVENT The MACHO collaboration (cf. IAUC 6068) reports the discovery of

a possible gravitational microlensing event in progress toward the Large Magellanic Cloud: “The star in question is located at α = 5^{h}17^{m}14^{s}.6, δ = −70^{o}46′59′′ (equinox 2000.0); it was constant at R = 19.8, V = 20.1 during 1993 and mid-1994, and it has brightened by 0.75 mag during the two weeks prior to Oct. 13, while its color appears unchanged. Magnitude estimates from the CCD on the 1.27-m telescope at Mt. Stromlo: 1994 Sept. 25.72 UT, V = 19.87; Oct. 5.65, 19.63; 10.70, 19.50 ; 13.65, 19.36 (errors typically 0.10 mag). If this event is due to gravitational microlensing, a preliminary fit predicts that the star should reach a peak of V < 19.2 between Oct. 15 and 30, and then decline symmetrically. As the light curve is currently incomplete and the star is faint, it should be emphasized that the microlensing interpretation is still tentative. Because this star is faint, the data obtained with our routine observations at the Mt. Stromlo 1.27-m telescope may not be sufficient for an accurate determination of the light curve shape. Thus, accurate photometry and low-resolution spectroscopy of this star at regular intervals during the next 14 days will be very valuable to test the microlensing interpretation. Light curves and a finding chart are available by anonymous FTP (address lensing.physics.ucsb.edu in the directory macho/alert/lmc1; this directory will be periodically updated with new light curves and microlensing fits as the data come in). Potential observers are requested to contact D. Bennett (telephone USA 510-423-0656; e-mail bennett@igpp.llnl.gov) to coordinate observations.”

OJ 287 AND 3C 66A M. Fiorucci and G. Tosti, Dipartimento di Fisica, Universitá di Pe-

rugia, write: “During our photometric monitoring program of blazars, we have found that OJ 287 and 3C 66A are in a high-luminosity stage, as indicated by the following photometry obtained with the 0.4-m reflector (+ CCD) of the Perugia Automatic Observatory using Cousins BVRI bands: OJ 287: May 15 UT, V = 15.50; Oct. 11.11, B = 14.86, V = 14.47, R = 14.04, I = 13.40; Oct. 12.16, 14.81, 14.35, 13.96, 13.34. 3C 66A: Oct. 11.07, B = 14.55, V = 14.23; Oct. 12.04, 14.52, 14.15.”

1994 October 12 Daniel W. E. Green